# Utilizing Contextual Clues and Role Correlations for Enhancing Document-level Event Argument Extraction

### v2

@@ 6 Conclusion
In this paper, we introduce the CARLG model for document-level event argument extraction (EAE). CARLG comprises two novel modules: Contextual Clues Aggregation (CCA) and Role-Based Latent Information Guidance (RLIG). CCA effectively aggregates non-argument clue words to capture beneficial contextual information, while RLIG provides guidance based on latent role representations that capture semantic correlations between some roles. Diverse experiments show that our CARLG model surpasses current SOTA EAE models on three large-scale document-level datasets.
While our CARLG model has exhibited remarkable performance in document-level event argument extraction (EAE) tasks with known event triggers, we acknowledge its limitation in depending on the availability of event triggers. In practical situations, event triggers may not always be supplied or known in advance. Therefore, in our future research, we intent to transfer our approach to encompass other EAE tasks that do not rely on explicit trigger words. We aim to conduct extensive experiments to assess its performance in such scenarios. Furthermore, we intend to explore the applicability of the CARLG framework in other information extraction tasks, such as relation extraction.

### v4

@@ 7 Conclusion
In this paper, we introduce the CARLG model for document-level event argument extraction. CARLG comprises two novel modules: Contextual Clues Aggregation (CCA) and Role-Based Latent Information Guidance (RLIG). CCA module effectively aggregates non-argument clue words to capture beneficial contextual information, while RLIG module provides guidance based on latent role representations that capture semantic correlations between some roles. Diverse experiments show that our CARLG model surpasses current SOTA EAE models on three large-scale document-level datasets.
While our CARLG model has exhibited remarkable performance in document-level event argument extraction tasks with known event triggers, we acknowledge its limitation in depending on the availability of event triggers. In practical situations, event triggers may not always be supplied or known in advance. Therefore, in our future research, we intent to transfer our approach to encompass other EAE tasks that do not rely on explicit trigger words. We aim to conduct extensive experiments to assess its performance in such scenarios. Furthermore, we intend to explore the applicability of the CARLG framework in other information extraction tasks, such as relation extraction.
[…]